Estimate the mass of air in this room.
1
Expert's answer
2020-12-28T08:42:31-0500

The molar mass of air is 28.97 g/mol.

For example, volume of room will be "40 \; m^3 \; (5 \; m \times 4 \; m \times 2 \; m)."

"V = 40 \; m^3 = 40000 \; L"

The temperature will be 20 ºC, the pressure 1 atm.

T = 20 + 273 = 293 K

p = 1 atm

Ideal gas law:

pV = nRT

R=0.082058 L×atm/(K×mol)

"n = \frac{pV}{RT} \\

n = \frac{1 \times 40000}{0.082058 \times 293} = 1663.68 \;mol \\

m_{air} = 1663.68 \times 28.97 \\

= 48196.8 \;g \\

= 48.19 \; kg"

Answer: 48.19 kg